PFI-ZEKE SPECTROSCOPY OF COPPER-AROMATIC NITROGEN HETEROCYCLIC COMPLEXES

\underline{XU WANG}, JASON F. FULLER, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY, 40605-0055.

Single-photon pulsed field ionization-zero electron kinetic energy (PFI-ZEKE) photoelectron spectroscopy has been used to study the complexes of copper with pyridine (C_{5}H_{5}N), 2,2’-bipyridine (C_{10}H_{8}N_{2}) and 1,10-phenanthroline (C_{12}H_{8}N_{2}). In all cases, adiabatic ionization potentials and Cu-ring vibrational frequencies were measured from the ZEKE spectra. Additional ring-based vibrations were observed for Cu-bipyridine and -phenanthroline. All three Cu complexes were determined to be in C_{2v} symmetry with the metal atom σ bonding to the nitrogen of pyridine or two nitrognes of the polypyridines. Since the most stable conformation of the free 2,2’-bipyridine ligand is coplanar with trans-standing nitrogens, Cu coordination has induced a ring rotation of 180^{o} to bring the two nitrogens in a cis configuration.